# FlexEdit: Marrying Free-Shape Masks to VLLM for Flexible Image Editing

### v1

@@ 4.1 Experiments Setting
Multiple Image: We include PNSR, CLIP-T, CLIP-I (Radford et al. 2021), and DINOv2 (Oquab et al. 2023) for the metric of foreground quality assessing, where CLIP-T measures how performance of the edited foreground aligns with the text label. PNSR, CLIP-I and DINOv2 assess the visual similarity between the generated images and the GT foreground images. LPIPS was introduced in the whole image evaluation while CLIP-T is excluded since CLIP-T label can not be annotated in the whole image setting.
Single Image: For single-image editing, in line with SmartEdit (Huang et al. 2024), we use PSNR, SSIM, LPIPS to evaluate background similarity, and CLIP-T to assess the alignment of the results’ foreground with the text label.
@@ 4.2 Comparisons with SOTA models
Multiple Image Editing
@@
Single Image Editing
@@ 4.3 Ablation study on MEA
We designed two different structures and compared their performance metrics in both single image and multiple image editing to validate the effectiveness of each module within the adapter. First, we remove the Cross Attention (CA) layer and directly used the output of the QFormer as the scene feature input into the following decouple cross attention, without the scene picture interaction. We also change the Decouple Cross Attention (DCA) structure with a general cross attention and examine the MEA effectiveness in image fusion. As shown in Table 3, in multiple image editing, the CA and DCA can help boost the model performance, both seperately and jointly. For single image editing, as shown in Table 4, though the MEA module does not work in this mode, it can learn during the joint training process, and CA, DCA involving shows a better result for the single image performance.
@@ 5 Conclusion
Appendix A Implementation Details
A.1 Training details
The LLaVAv1.1-7B (Liu et al. 2024) and QFormer (Li et al. 2023c) pretrained by SmartEdit (Huang et al. 2024) and the Stable Diffusion-1.5 (Rombach et al. 2022b) are the VLLM and diffusion model pretrained weights. The base LLM for LLaVA1.1-7B is Vicuna-7B  (Zheng et al. 2024). FlexEdit is trained for 12000 steps on 4 NVIDIA A100 GPUs with batch size 4 and gradient accumulation 8, using AdamW (Loshchilov and Hutter 2017) optimizer with $1\times 10^{-5}$ learing rate. Number $N$ of image new tokens is set to 32.
For the fine-tuning process, we employed DeepSpeed  (Aminabadi et al. 2022) Zero-2 to perform LoRA (Hu et al. 2021) fine-tuning with the following parameters: LoRA Rank = 8, LoRA Alpha = 16, and LoRA Dropout = 0.05. The base model is frozen and the LoRA fine-tuning target modules are query project layers and value project layers.
A.2 Data Construction Pipeline
In this section, we present detail methodology of the FSMI-Edit benchmark and the benchmark data construction pipeline.
Figure 8: In the multiple images editing datasets creation process, we use Adobe Photoshop to extract the full mask from the result image, then redraw the mask in different shape, and write prompts based on the task to serve as input instructions.
Figure 9: In the single image editing datasets creation process, we use Adobe Photoshop to draw different shaped masks in specified areas of the scene image [pic1] based on the task requirements. Then we write prompts based on the task to serve as input instructions.
To enable FlexEdit to handle complex scenes, intricate instructions, and free-shape mask editing, we have designed a data construction pipeline to build the self-construction dataset used in training and FSMI-Edit benchmark. These dataset and benchmark primarily include everyday items, indoor objects, and outdoor scenes. We have carefully designed eight types of free-shape masks to simulate user input preferences (e.g. casually and unprecisely), allowing for a more accurate evaluation of free-shape mask-based image editing performance.
The images in the benchmark come from real-world photography and internet sources. Images captured in real life are cropped to a suitable size to meet the benchmark requirements. Images collected from the internet have variations in saturation, contrast, and other parameters, which could affect the image editing results. To rectify this, we use Adobe Photoshop [^1] https://www.photoshop.com/ to process these images and obtain more harmonious images. Additionally, Adobe Photoshop integrates a segmentation model, allowing for direct segmentation operations during image processing. Combined with Adobe Photoshop’s AI filling technology, it can process images more conveniently and improve work efficiency. Therefore, we use Adobe Photoshop to prepare the benchmark.
In the benchmark, we mark the captured or collected scene images as [pic1], the object images as [pic2], the prompts written according to task requirements as prompt, and the various drawn shapes of masks as mask. The creation of the multiple images editing datasets is divided into two parts: images captured in real-life scenarios and images collected from the internet.
For images captured in real-life, tasks are completed by moving, placing, etc. to accomplish task requirements.Then we use Adobe Photoshop to extract the complete mask of specific objects from the images. On these complete masks, we draw eight free-shape masks that conform to user input habits, such as circles, rectangles, triangles, irregular shapes, and their hollow versions. Finally, we write prompts based on the task requirements to complete the dataset creation.
For images collected from the internet, The images collected on the Internet take the dog and polar bear in Figure 1 as an example. We use Adobe Photoshop to remove the dog on the right side of the scene image [pic1] and fill in the background to obtain the final scene image. Subsequently, we use the integrated segmentation model to extract the polar bear from the object image [pic2], then process the polar bear and fuse it into the scene image for further optimization. The extracted mask is drawn to obtain masks of different shapes. Finally, prompts are written according to the task description. The creation of the multiple images editing benchmark are completed.
Similarly, in the single image editing benchmark, we use Adobe Photoshop to draw free-shape masks in specified locations according to the task requirements, and prompts are written based on the task.
The data production pipeline is illustrated in Figure 8 and Figure 9. Figure 1 shows the production pipeline of the multiple images editing benchmark, and Figure 2 shows the production pipeline of the single image editing benchmark. The images of our benchmark come from images taken in the real world (such as the blue doll in Figure 1) and images collected from the Internet (such as the polar bear and dogs in Figure 1). Our benchmark tasks are divided into two categories: Add and Replace. For example, in Figure 1, the tasks include: “Add the blue doll from [pic2] to [pic1], with the white mask serving as a positional guide” and “Replace the dog in the white mask in [pic1] with the polar bear in [pic2]”. In Figure 2, the tasks include: “Add a yellow lemon in the white mask of [pic1]” and “Replace the fruit in the white area of [pic1] with a tomato”.
A.3 Dataset Composition
Figure 10: This pie chart illustrates the input distribution of datasets during the training process. The first chart shows the proportion of four different datasets used in training, while the second chart indicates the proportion of single-image versus multi-image inputs within the image editing dataset.
During the training process, we employed a mixed training approach, using four different datasets as input with specific probability ratios. As shown in Figure 10, the segmentation datasets modified from COCO and GRefCOCO account for 15%, the VQA dataset Mimic-it (Li et al. 2023a) accounts for 10%, the image editing datasets adapted from MagicBrush (Zhang et al. 2024), ReasonEdit (Huang et al. 2024), and STRAT (Zhang et al. 2022) account for 65%, and our self-constructed image editing dataset account for 10%. Within the whole image editing datasets, single-image editing tasks made up 20%, while multi-image editing tasks comprised 80%.
Appendix B Qualitative comparisons
Figure 11: Multiple images editing comparison on circle mask, circle open mask, irregular mask, irregular open mask.
Figure 12: Multiple images editing comparison on rectangle mask, rectangle open mask, triangle mask, triangle open mask.
Figure 13: Single image editing comparison on circle mask, circle open mask, irregular mask, irregular open mask.
Figure 14: Single images editing comparison on rectangle mask, rectangle open mask, triangle mask, triangle open mask.
In this section, we provide a qualitative comparison between FlexEdit with other method on 8 different types in FSMI-Edit Benchmark. The multiple images comparisons are shown in Figure 11 and Figure 12. The single image comparisons are shown in Figure 13 and Figure 14.
There are two situations for multiple models on the benchmark: one where the mask closely matches with the subject, and another where the mask and subject are less well-matched.
Some of the compared models either alter the background image (as seen with InstructAny2Pix) or change the size of the object (as seen with Anydoor and Largen), leading to a decrease in visual quality. For example, in the first row of Figure 5, the user’s task is to add the bag in [pic2] to the white mask in [pic1], Anydoor and Largen generate the bag within the masked area, resulting in a bag size that is inconsistent with the user’s specifications. InstructAny2Pix, on the other hand, alters the overall image style. In contrast, our FlexEdit integrates the specified bag into the designated position in the scene with an appropriate size. The visual quality of the images generated by FlexEdit is superior to that of the other models.
If the mask does not match well with the subject, particularly when there are hollows in the mask shape, the compared models either fail to interpret the free-shape mask correctly (such as InstructPix2Pix), alter the background image, or only generate images within the masked area (as seen with Anydoor Largen). Specifically, in the third row of Figure 5, the user’s task is to replace the basketball in [pic1] with the bag in [pic2], Anydoor generates the bag within the mask area, Largen fails to understand the free-shape mask, and InstructAny2Pix alters the overall image style. In contrast, our FlexEdit model preserves the background and replaces the basketball with the bag in the scene, maintaining the correct size and positioning. This further highlights the strong free-shape mask interpretation and image editing capabilities of our FlexEdit model.

### v2

@@ 4.1.2 Metrics
Multiple Image: We include PNSR, CLIP-T, CLIP-I [24], and DINOv2 [20] for the metric of foreground quality assessing, where CLIP-T measures how performance of the edited foreground aligns with the text label. PNSR, CLIP-I and DINOv2 assess the visual similarity between the generated images and the GT foreground images. LPIPS was introduced in the whole image evaluation while CLIP-T is excluded since CLIP-T label can not be annotated in the whole image setting.
Single Image: For single-image editing, in line with SmartEdit [9], we use PSNR, SSIM, LPIPS to evaluate background similarity, and CLIP-T to assess the alignment of the results’ foreground with the text label.
@@ 4.2.1 Baselines
4.2.2 Multiple Image Editing
@@
4.2.3 Single Image Editing
@@
Figure 7: Qualitative results comparison on single image edit.
@@
Table 2: Comparison in single image editing domain.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CLIP-T $\uparrow$ InstructPix2Pix 19.9414 0.8016 0.1043 20.7010 InstructAny2Pix 15.5149 0.5919 0.2751 20.3758 SmartEdit 21.9535 0.7620 0.0781 20.4910 FlexEdit 22.8081 0.7712 0.0727 20.7081
@@ 4.3 Ablation study on MEA
To validate our contribution, the ablation study was intentionally focused on the MEA’s key components: the Cross Attention (CA) and Decouple Cross Attention (DCA) modules. This follows the standard experimental practice of isolating variables to clearly demonstrate the effectiveness of a new component. We designed two variant structures and compared their performance metrics in both single and multiple image editing to validate the effectiveness of each module. First, we remove the Cross Attention (CA) layer and directly used the output of the QFormer as the scene feature input into the following decouple cross attention, without the scene picture interaction. We also change the Decouple Cross Attention (DCA) structure with a general cross attention and examine the MEA effectiveness in image fusion. As shown in Table 3, in multiple image editing, the CA and DCA can help boost the model performance, both seperately and jointly. For single image editing, as shown in Table 4, though the MEA module does not work in this mode, it can learn during the joint training process, and CA, DCA involving shows a better result for the single image performance.
4.4 Limitations and Failure Case Analysis
Despite the strong performance of FlexEdit, our work has some limitations and potential areas for future improvement.
Choice of VLLM Backbone. We chose LLaVA to ensure a controlled and fair comparison with SOTA methods like SmartEdit, allowing us to isolate the benefits of our proposed contributions. While this provides a clear comparison, future work could explore performance improvements with alternative or more recent VLLM backbones, such as Qwen-VL[1].
Failure Cases. We have identified three primary modes of failure:
  • VLLM Grounding Failure: When a free-shape mask is highly ambiguous (e.g., covering two distinct objects simultaneously), the VLLM may misinterpret the user’s intent, leading to an incorrect edit.
  • Semantic Misinterpretation: In some cases, the model may generate an object that is semantically related to the instruction but is contextually incorrect (e.g., generating a tiger when a cat was requested).
  • Background Artifacts: Minor, unintended changes to the background style or texture can occasionally occur. This is a known challenge in end-to-end diffusion generation frameworks that can affect the overall image quality.